Consider the following system of linear algebraic equations 𝐴𝑥=𝑏:


[1 1 3 [ X = [2


5 3 1 Y 3


2 3 1] Z] -1]


Solve the system using (you can compare your solution with MATLAB)


1. Cramer’s Rule


2. Gauss elimination without pivoting


3. Gauss-Jordan elimination


4. LU factorization


1
Expert's answer
2022-02-21T16:24:28-0500

"A=\begin{pmatrix}
 1 & 1&3 \\
 5&3&1\\
2&3&1
\end{pmatrix}, x=\begin{pmatrix}
 x \\
 y\\z
\end{pmatrix},b=\begin{pmatrix}
 2 \\
 3\\-1
\end{pmatrix}"

1. Cramer’s Rule

"\Delta=\begin{vmatrix}
 1 & 1&3 \\
 5&3&1\\
2&3&1
\end{vmatrix}=3+2+45-18-5-3=24\\
\Delta_x=\begin{vmatrix}
 2 & 1&3 \\
 3&3&1\\
-1&3&1
\end{vmatrix}=6-1+27+9-3-6=32\\
\Delta_y=\begin{vmatrix}
 1 & 2&3 \\
 5&3&1\\
2&-1&1
\end{vmatrix}=3+4-15-18-10+1=-35"

"\Delta_z=\begin{vmatrix}
 1 & 1&2 \\
 5&3&3\\
2&3&-1
\end{vmatrix}=-3+6+30-12+5-9=17"

"x=\frac{\Delta _1}{\Delta}=\frac{32}{24}=\frac{4}{3}\\
y=\frac{\Delta _2}{\Delta}=\frac{-35}{24}=-\frac{35}{24}\\
z=\frac{\Delta _3}{\Delta}=\frac{17}{24}"

2. Gauss elimination without pivoting

"\begin{pmatrix}
 1 & 1&3 &|&2\\
 5&3&1&|&3\\
2&3&1&|&-1
\end{pmatrix}\to"

2r+1r(-5)

3r+1r(-2)

"\begin{pmatrix}
 1 & 1&3 &|&2\\
 0&-2&-14&|&-7\\
0&1&-5&|&-5
\end{pmatrix}\to"

2r"\leftrightarrow" 3r

"\begin{pmatrix}
 1 & 1&3 &|&2\\
0&1&-5&|&-5\\
 0&0&-24&|&-17\\
\end{pmatrix}"

"x+y+3z=2\\
-2y-14z=-7\\
-24z=-17\mapsto z=\frac{17}{24}"

"-2y-14\cdot\frac{17}{24}=-7\mapsto x_2=-\frac{35}{24}\\
x-\frac{35}{24}+3\cdot\frac{17}{24}=2\mapsto x=\frac{3}{4}"


3. Gauss-Jordan elimination

"\begin{pmatrix}
 1 & 1&3 &|&2\\
 5&3&1&|&3\\
2&3&1&|&-1
\end{pmatrix}\to"

2r+1r(-5)

3r+1r(-2)

"\begin{pmatrix}
 1 & 1&3 &|&2\\
 0&-2&-14&|&-7\\
0&1&-5&|&-5
\end{pmatrix}\to"

2r"\leftrightarrow" 3r

"\begin{pmatrix}
 1 & 1&3 &|&2\\
0&1&-5&|&-5\\
 0&-2&-14&|&-7
\end{pmatrix}\to"

3r+2r(2)

"\begin{pmatrix}
 1 & 1&3 &|&2\\
0&1&-5&|&-5\\
 0&0&-24&|&-17\\

\end{pmatrix}\to"

3r("\frac{-1}{24}" )

"\begin{pmatrix}
 1 & 1&3 &|&2\\
 0&1&-5&|&-5\\
0&0&1&|&\frac{17}{24}
\end{pmatrix}\to"

2r+3r(5)

1r+3r(-3)

"\begin{pmatrix}
 1 & 1&0 &|&-\frac{3}{24}\\
 0&1&0&|&-\frac{35}{24}\\
0&0&1&|&\frac{17}{24}
\end{pmatrix}\to"

1r+2r(-1)

"\begin{pmatrix}
 1 & 0&0 &|&\frac{32}{24}\\
 0&1&0&|&-\frac{35}{24}\\
0&0&1&|&\frac{17}{24}
\end{pmatrix}"

"x=\frac{32}{24}=\frac{4}{3}, y=-\frac{35}{24}, z=\frac{17}{24}"


4. LU factorization

"L=\begin{pmatrix}
 1 & 0&0 \\
 5 & 1&0\\
2&-\frac{1}{2}&1
\end{pmatrix}"

5: 2:

"\begin{pmatrix}
 1 & 1&3 \\
 0&-2&-14\\
0&1&-5
\end{pmatrix}"

"-\frac{1}{2}" :

"\begin{pmatrix}
 1 & 1&3 \\
 0&-2&-14\\
0&0&-12
\end{pmatrix}=U"

"LY=B\\"

"\begin{pmatrix}
 1 & 0&0 \\
 5 & 1&0\\
2&-\frac{1}{2}&1
\end{pmatrix}\begin{pmatrix}
 y_1\\
 y_2\\
y_3
\end{pmatrix}=\begin{pmatrix}
 2\\
 3\\
-1
\end{pmatrix}"

"y_1=2\\
5y_1+y_2=3\mapsto y_2=-7\\
2y_1-\frac{1}{2}y_2+y_3=-1\mapsto y_3=-\frac{17}{2}"

"UX=Y\\"

"\begin{pmatrix}
 1 & 1&3 \\
 0&-2&-14\\
0&0&-12
\end{pmatrix}\begin{pmatrix}
 x \\
 y\\z
\end{pmatrix}=\begin{pmatrix}
 2 \\
 -7\\-\frac{17}{2}
\end{pmatrix}"

"x+y+3z=2\mapsto x=-\frac{32}{24} \\
-2y-14z=-7\mapsto y=-\frac{35}{24}\\
-12z=-\frac{17}{2}\mapsto z=\frac{17}{24}"